OpenStudy (loser66):

If f(x) is differentiable at x0. Prove it continuous at x0 Please, help

OpenStudy (loser66):

@xapproachesinfinity hey kid, help old man please

OpenStudy (xapproachesinfinity):

hey old man long time

OpenStudy (xapproachesinfinity):

i think theorem state if f is differentiable applies f is continuous so you are proving the theorem

OpenStudy (xapproachesinfinity):

so let see f differentiable at x0 means the limits exist f'(x0)

OpenStudy (xapproachesinfinity):

so \(\lim_{x\to x_0} \frac{f(x)-f(x_0)}{x-x_0}=\text{exist}\)

OpenStudy (xapproachesinfinity):

you have to work from here

OpenStudy (xapproachesinfinity):

first let's state the definition of f being cont at x0 that is to say \(\lim_{x\to x_0} f(x)=f(x_0)\)

OpenStudy (loser66):

Sorry, kid, old man was kicked out of the site.

OpenStudy (loser66):

But I think you got circulation argument: differentiable --> continuous --

OpenStudy (loser66):

However, I got it. \(f(x) - f(x_0) = \dfrac{f(x) -f(x_0)}{x-x_0}(x-x_0)\) \(lim_{x\rightarrow x_0} |f(x) - f(x_0)| = lim_{x\rightarrow x_0}\dfrac{|f(x) -f(x_0)}{|x-x_0|}|x-x_0|\) since f(x) differentiable at \(x_0\), the limit exists and as x approaches \(x_0\) \(|x-x_0|=0\) That gives us \(lim_{x\rightarrow x_0}|f(x) -f(x_0)|=0\) or \(lim_{x\rightarrow x_0}f(x) = f(x_0)\) That shows f(x) continuous at x0

OpenStudy (xapproachesinfinity):

there is no circulation it is one way f different implies f continuous so you start with f different stating the definition limf(x)-f(x0)/x-x0=f'(x0) multilply both sides by (x-x0) we get lim(f(x)-f(x0)=f'(x0)(x-x0) as x tends to x0 the quanties (x-x0) shrinks to zero (limit concept) then we are safe to say lim(f(x)-f(x0)=f'(x0)0=0 therefore lim(f(x))=f(x0) hence f is continuous at x0

OpenStudy (xapproachesinfinity):

a subtlety here is the |x-c| is zero but it is not zero however we write as it is zero